OpenStudy (anonymous):

Find the slope of the line containing the pair of points (8,-1) & (-8,-1).

OpenStudy (anonymous):

You do rise over run, \[m=(y2-y1)\div(x2-x1)\] so m=(1-(-1))/(-8-8)..m=0/-16, the slope is zero, if you draw the graph you will see that it is a horizontal line

OpenStudy (anonymous):

*that's not right, i didnt see the -1

OpenStudy (anonymous):

scratch that, lol, the equation is right, m=0, I just put 1 as y2 instead of -1 in the slope equation

OpenStudy (anonymous):

Ok so what is the slope of the line?

OpenStudy (anonymous):

the slope (m)=0/16, which is a slope of 0....m=0

OpenStudy (anonymous):

when in doubt draw a quick graph

OpenStudy (anonymous):

Ok thank you. Sorry it took so long to reply. Was writing a narrative essay, fun fun. NOT.

OpenStudy (anonymous):

haha, np, essays are horrible, well good luck